Three charges are placed at the vertices of an equilateral triangle. For what value of Q both sign and magnitude does the total electric field vanish at the center of the triangle? 


1
Expert's answer
2021-07-01T03:59:02-0400

Electric field at "q_1=q" charge

"E_1=\frac{kq}{(\frac{a}{\sqrt3})^2}=3\frac{kq}{a^2}"

"E_2=\frac{kQ}{(\frac{a}{\sqrt3})^2}=3\frac{kQ}{a^2}"

"E_3=\frac{kQ}{(\frac{a}{\sqrt3})^2}=3\frac{kQ}{a^2}"

E2andE3 resultant

"E'=\sqrt{E_1^2+E_2^2+2E_1E_2cos\phi}"

"\phi=60°"

"E_1=E_2=E"

"cos60°=\frac{1}{2}"

"E'=\sqrt{2}E"

"E'=\sqrt2\frac{kQ}{a^2}"

Tringle Centre all three charge electric field zero

When

"E'=E_1"

"3\frac{kq}{a^2}=3\sqrt{2}\frac{kQ}{a^2}"

"Q=\frac{q}{\sqrt2}"